OpenStudy (anonymous):

What are the x- and y-intercepts of the graph of y = 2x2 + 4x - 6? a. x-intercepts: (1, 0) and (-3, 0) y-intercept: (0, -3) b. x-intercepts: (1, 0) and (-3, 0) y-intercept: (0, -6) c. x-intercepts: (-1, 0) and (3, 0) y-intercept: (0, -3) d. x-intercepts: (-1, 0) and (3, 0) y-intercept: (0, -6)

OpenStudy (anonymous):

Plug in 0 in y to get the x intercept, and then plug in 0 in x to get the y intercept.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

y=2(0)^2 +4(0) - 6 y= 0 +0-6 y=-6 (0,-6) and 0=2x^2 +4x -6 0= 2x^2-2x +6x -6 0= 2x(x-1)+6(x-1) 0= (2x+6)(x-1) x=-3 or x=1 so, (-3,0) or (1,0)